# ReadMe++: Benchmarking Multilingual Language Models for Multi-Domain Readability Assessment

### v3

@@ 2 Related Work
Examples of Data Sources — Full list for all languages in Appendix A Domain (Abrv) # Arabic (ar) English (en) Hindi (hi) Captions (Cap) 9 Images ElJundi et al. (2020) Videos Wang et al. (2019) Movies Lison and Tiedemann (2016) Dialogue (Dia) 7 Open-domain Naous et al. (2020) Negotiation He et al. (2018) Task-oriented Malviya et al. (2021) Dictionaries (Dic) 2 Dictionaries (almaany.com) Dictionaries (dictionary.com) Entertainment (Ent) 4 Jokes (almrsal.com) Jokes Weller and Seppi (2019) Jokes (123hindijokes.com) Finance (Fin) 3 Finance Malo et al. (2014) Forums (For) 7 QA Websites (hi.quora.com) StackOverflow Tabassum et al. (2020) Reddit (reddit.com) Guides (Gui) 6 Online Tutorials (ar.wikihow.com) Code Documentation (mathworks.com) Cooking Recipes (narendramodi.in) Legal (Leg) 9 UN Parliament Ziemski et al. (2016) Constitutions (constitutioncenter.org) Judicial Rulings Kapoor et al. (2022) Letters (Let) 3 Letters (oflosttime.com) Literature (Lit) 3 Novels (hindawi.org/books/) History (gutenberg.org) Biographies (Public Domain Books) Medical Text (Med) 1 Clinical Reports Uzuner et al. (2011) News Articles (New) 2 Sports Alfonse and Gawich (2022) Economy Misra (2022) Poetry (Poe) 5 Poetry (aldiwan.net) Poetry (poetryfoundation.org) Poetry (hindionlinejankari.com) Policies (Pol) 7 Olympic Rules (specialolympics.org) Contracts (honeybook.com) Code of Conduct (lonza.com) Research (Res) 15 Politics (jcopolicy.uobaghdad.edu.iq) Science & Engineering (arxiv.org) Economics (journal.ijarms.org) Social Media (Soc) 3 Twitter Zheng et al. (2022) Twitter Zheng et al. (2022) Twitter Zheng et al. (2022) Speech (Spe) 4 Public Speech (state.gov/translations) Public Speech (whitehouse.gov) Ted Talks (ted.com/talks) Statements (Sta) 6 Quotes (arabic-quotes.com) Rumours Zheng et al. (2022) Quotes (wahh.in) Textbooks (Tex) 3 Business (hindawi.org/books/) Agriculture (open.umn.edu) Psychology (ncert.nic.in) User Reviews (Rev) 12 Products ElSahar and El-Beltagy (2015) Books (goodreads.com) Movies (hindi.webdunia.com) Wikipedia (Wik) 1 Wikipedia (wikipedia.com) Wikipedia (wikipedia.com) Wikipedia (wikipedia.com) Total 112
@@ 4.1 Supervised & Prompting Methods
We fine-tune LMs to classify sentence readability. We compare multilingual models, mBERT Devlin et al. (2019) and XLM-R Conneau et al. (2020), to monolingual models that include BERT Devlin et al. (2019) for English, AraBERT Antoun et al. (2020) and ArBERT Abdul-Mageed et al. (2021) for Arabic, CamemBERT for French Martin et al. (2020), and RuBERT Kuratov and Arkhipov (2019) for Russian. For Hindi, we use MuRIL Khanuja et al. (2021) and IndicBERTv2 Kakwani et al. (2020), both pre-trained on 12 Indian languages. We also consider encoder-decoder LMs, mT5 Xue et al. (2021), Aya Üstün et al. (2024), and AraT5 Elmadany et al. (2022). We fine-tune LMs for 20 epochs using the cross-entropy loss and the Adam optimizer and tune the learning rate in the set $\{1e^{-5},1e^{-6},1e^{-7}\}$. We select checkpoints based on the best performance on the validation set. We report the average of 5 runs with different random initialization seeds.
@@
We perform in-context learning using GPT3.5, GPT4, and Llama2-7B Touvron et al. (2023). We provide the LMs with a definition of readability and the descriptors of the six CEFR levels provided to our human annotators. We show the model five randomly sampled in-context examples from the train set and their corresponding CEFR levels, then ask the model to assess the readability of a new sentence based on the CEFR scale. Prompt details can be found in Appendix D.3.
@@
Figure 4: Pearson correlation ($\rho$) of fine-tuned multilingual and monolingual LMs, as well as prompted GPT3.5, GPT4 and Llama2 models with 5-shots, on the test set of ReadMe++. The small (S), base (B), and large (L) sizes of the models are used. We report the min/max/average of performance across 5 runs using random seeds for fine-tuning initialization, or random sets of demonstrations in prompting.
@@ 4.1.1 Results
Fine-tuned models were able to achieve high correlation levels in the 0.7-0.9 range, with larger models showing improved performance. Overall, mT5L was among the best-performing fine-tuned models across all languages. However, the performance of GPT3.5, GPT4, and Llama2 with 5-shot examples was lower than that of fine-tuned models in all languages.
@@ 4.2.1 Results
Table 5: Zero-shot cross-lingual transfer results using XLMRL. Models fine-tuned on English data (en) of ReadMe++ significantly outperform models fine-tuned with CEFR-SP Arase et al. (2022) or CompDS Brunato et al. (2018) in transfer to Arabic (ar), Hindi (hi), French (fr), Russian (ru), Italian (it), and German (de).
@@ Acknowledgments
The authors would like to thank Nour Allah El Senary, Govind Ramesh, Suraj Mehrotra, and Ryan Punamiya for their help in data annotation. This research is supported in part by the NSF awards IIS-2144493 and IIS-2112633, ODNI and IARPA via the HIATUS program (contract 2022-22072200004). The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of NSF, ODNI, IARPA, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
@@ Appendix C Traditional Metrics
Pre-training Sources Model #Params Wiki News Books CC Multilingual LMs mBERT 177M $\checkmark$ XLMRB 278M $\checkmark$ XLMRL 559M $\checkmark$ mT5S 60M $\checkmark$ mT5B 220M $\checkmark$ mT5L 770M $\checkmark$ Aya 13B $\checkmark$ Monolingual Arabic LMs AraBERTB 135M $\checkmark$ $\checkmark$ AraBERTL 369M $\checkmark$ $\checkmark$ $\checkmark$ ArBERT 163M $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ AraT5B 220M $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Monolingual French LMs CamemBERTB 110M $\checkmark$ CamemBERTL 335M $\checkmark$ Monolingual English LMs BERTB 110M $\checkmark$ $\checkmark$ BERTL 350M $\checkmark$ $\checkmark$ Indian LMs MuRILB 237M $\checkmark$ $\checkmark$ MuRILL 506M $\checkmark$ $\checkmark$ IndicBERTv2B 278M $\checkmark$ $\checkmark$ Monolingual Russian LMs RuBERTB 180M $\checkmark$
@@ D.3 Few-shot Prompt
Rate the following sentence on it’s readability level. The readabilty is defined as the cognitive load required to understand the meaning of the sentence. Rate the readabilty on a scale from very easy to very hard. Base your scores off the CEFR scale for L2 Learners. You should use the following key: 1 = Can understand very short, simple texts a single phrase at a time, picking up familiar names, words and basic phrases and rereading as required. 2 = Can understand short, simple texts on familiar matters of a concrete type 3 = Can read straightforward factual texts on subjects related to his/her field and interest with a satisfactory level of comprehension. 4 = Can read with a large degree of independence, adapting style and speed of reading to different texts and purpose 5 = Can understand in detail lengthy, complex texts, whether or not they relate to his/her own area of speciality, provided he/she can reread difficult sections. 6 = Can understand and interpret critically virtually all forms of the written language including abstract, structurally complex, or highly colloquial literary and non-literary writings. EXAMPLES: Sentence: "[EX 1]" Given the above key, the readabilty of the sentence is (scale=1-6): [EX RATING 1] Sentence: "[EX 2]" Given the above key, the readabilty of the sentence is (scale=1-6): [EX RATING 2] ... Sentence: "[EX N]" Given the above key, the readabilty of the sentence is (scale=1-6): [EX RATING N] Sentence: "[SENTENCE]" Given the above key, the readability of the sentence is (scale=1-6):
Table 10: Prompt provided to GPT4, GPT3.5, and Llama2-7B models to assess in-context learning readability assessment capabilities.

### v4

@@ 2 Related Work
Examples of Data Sources — Full list for all languages in Appendix A Domain (Abrv) # Arabic (ar) English (en) Hindi (hi) Captions (Cap) 9 Images ElJundi et al. (2020) Videos Wang et al. (2019) Movies Lison and Tiedemann (2016) Dialogue (Dia) 7 Open-domain Naous et al. (2020) Negotiation He et al. (2018) Task-oriented Malviya et al. (2021) Dictionaries (Dic) 2 Dictionaries (almaany.com) Dictionaries (dictionary.com) Entertainment (Ent) 4 Jokes (almrsal.com) Jokes Weller and Seppi (2019) Jokes (123hindijokes.com) Finance (Fin) 3 Finance Malo et al. (2014) Forums (For) 7 QA Websites Nakov et al. (2016) StackOverflow Tabassum et al. (2020) Reddit (reddit.com) Guides (Gui) 6 Online Tutorials (ar.wikihow.com) Code Documentation (mathworks.com) Cooking Recipes (narendramodi.in) Legal (Leg) 9 UN Parliament Ziemski et al. (2016) Constitutions (constitutioncenter.org) Judicial Rulings Kapoor et al. (2022) Letters (Let) 3 Letters (oflosttime.com) Literature (Lit) 3 Novels (hindawi.org/books/) History (gutenberg.org) Biographies (Public Domain Books) Medical Text (Med) 1 Clinical Reports Uzuner et al. (2011) News Articles (New) 2 Sports Alfonse and Gawich (2022) Economy Misra (2022) Poetry (Poe) 5 Poetry (aldiwan.net) Poetry (poetryfoundation.org) Poetry (hindionlinejankari.com) Policies (Pol) 7 Olympic Rules (specialolympics.org) Contracts (honeybook.com) Code of Conduct (lonza.com) Research (Res) 15 Politics (jcopolicy.uobaghdad.edu.iq) Science & Engineering (arxiv.org) Economics (journal.ijarms.org) Social Media (Soc) 3 Twitter Zheng et al. (2022) Twitter Zheng et al. (2022) Twitter Zheng et al. (2022) Speech (Spe) 4 Public Speech (state.gov/translations) Public Speech (whitehouse.gov) Ted Talks (ted.com/talks) Statements (Sta) 6 Quotes (arabic-quotes.com) Rumours Zheng et al. (2022) Quotes (wahh.in) Textbooks (Tex) 3 Business (hindawi.org/books/) Agriculture (open.umn.edu) Psychology (ncert.nic.in) User Reviews (Rev) 12 Products ElSahar and El-Beltagy (2015) Books (goodreads.com) Movies (hindi.webdunia.com) Wikipedia (Wik) 1 Wikipedia (wikipedia.com) Wikipedia (wikipedia.com) Wikipedia (wikipedia.com) Total 112
@@ 4.1 Supervised & Prompting Methods
We fine-tune LMs to classify sentence readability. We compare multilingual models, mBERT Devlin et al. (2019) and XLM-R Conneau et al. (2020), to monolingual models that include BERT Devlin et al. (2019) for English, AraBERT Antoun et al. (2020) and ArBERT Abdul-Mageed et al. (2021) for Arabic, CamemBERT for French Martin et al. (2020), and RuBERT Kuratov and Arkhipov (2019) for Russian. For Hindi, we use MuRIL Khanuja et al. (2021) and IndicBERTv2 Kakwani et al. (2020), both pre-trained on 12 Indian languages. We also consider encoder-decoder LMs, mT5 Xue et al. (2021), Aya101 Üstün et al. (2024), and AraT5 Elmadany et al. (2022). We fine-tune for 20 epochs using the cross-entropy loss and the Adam optimizer and tune the learning rate in the set $\{1e^{-5},1e^{-6},1e^{-7}\}$. We select checkpoints based on the best performance on the validation set. We report the average of 5 runs with different random initialization seeds.
@@
We perform in-context learning using GPT3.5, GPT4 (Apr 2024), Llama2-7b Touvron et al. (2023), Llama3.1-8b Dubey et al. (2024), and Aya23-8b Aryabumi et al. (2024). We provide LMs with a definition of readability and the descriptors of the six CEFR levels. We show the model five randomly sampled in-context examples from the train set and their corresponding CEFR levels, then ask the model to assess the readability of a new sentence based on the CEFR scale. Prompt details can be found in Appendix D.3.
@@
Figure 4: Pearson correlation ($\rho$) of fine-tuned multilingual and monolingual LMs, as well as prompted GPT3.5, GPT4, Aya23-8b, Llama2-7b, and Llama3.1-8b models with 5-shot examples, on the test set of ReadMe++. The small (S), base (B), and large (L) sizes of the models are used. We report the min/max/average of performance across 5 runs using random seeds for fine-tuning initialization, or random sets of demonstrations in prompting.
@@ 4.1.1 Results
Fine-tuned models were able to achieve high correlation levels in the 0.7-0.9 range, with larger models showing improved performance. Overall, mT5L was among the best-performing fine-tuned models across all languages. However, the performance of prompted causal models with 5-shot examples was lower than that of fine-tuned models in all languages.
@@ 4.2.1 Results
Table 5: Zero-shot cross-lingual transfer results using XLMRL. LMs fine-tuned on English data (en) of ReadMe++ significantly outperform LMs fine-tuned with CEFR-SP Arase et al. (2022) or CompDS Brunato et al. (2018) in transfer to Arabic (ar), Hindi (hi), French (fr), Russian (ru), Italian (it), and German (de).
@@ Acknowledgments
The authors would like to thank Nour Allah El Senary, Govind Ramesh, Suraj Mehrotra, and Ryan Punamiya for their help in data annotation. This research is supported in part by the NSF awards IIS-2144493 and IIS-2112633, NIH award R01LM014600, ODNI and IARPA via the HIATUS program (contract 2022-22072200004). The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of NSF, NIH, ODNI, IARPA, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
@@ Appendix C Traditional Metrics
Pre-training Sources Model #Params Wiki News Books CC Multilingual LMs mBERT 177M $\checkmark$ XLMRB 278M $\checkmark$ XLMRL 559M $\checkmark$ mT5S 60M $\checkmark$ mT5B 220M $\checkmark$ mT5L 770M $\checkmark$ Aya101 13B $\checkmark$ Monolingual Arabic LMs AraBERTB 135M $\checkmark$ $\checkmark$ AraBERTL 369M $\checkmark$ $\checkmark$ $\checkmark$ ArBERT 163M $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ AraT5B 220M $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ Monolingual French LMs CamemBERTB 110M $\checkmark$ CamemBERTL 335M $\checkmark$ Monolingual English LMs BERTB 110M $\checkmark$ $\checkmark$ BERTL 350M $\checkmark$ $\checkmark$ Indian LMs MuRILB 237M $\checkmark$ $\checkmark$ MuRILL 506M $\checkmark$ $\checkmark$ IndicBERTv2B 278M $\checkmark$ $\checkmark$ Monolingual Russian LMs RuBERTB 180M $\checkmark$
@@ D.3 Few-shot Prompt
Rate the following sentence on it’s readability level. The readabilty is defined as the cognitive load required to understand the meaning of the sentence. Rate the readabilty on a scale from very easy to very hard. Base your scores off the CEFR scale for L2 Learners. You should use the following key: 1 = Can understand very short, simple texts a single phrase at a time, picking up familiar names, words and basic phrases and rereading as required. 2 = Can understand short, simple texts on familiar matters of a concrete type 3 = Can read straightforward factual texts on subjects related to his/her field and interest with a satisfactory level of comprehension. 4 = Can read with a large degree of independence, adapting style and speed of reading to different texts and purpose 5 = Can understand in detail lengthy, complex texts, whether or not they relate to his/her own area of speciality, provided he/she can reread difficult sections. 6 = Can understand and interpret critically virtually all forms of the written language including abstract, structurally complex, or highly colloquial literary and non-literary writings. EXAMPLES: Sentence: "[EX 1]" Given the above key, the readability of the sentence is (scale=1-6): [EX RATING 1] Sentence: "[EX 2]" Given the above key, the readability of the sentence is (scale=1-6): [EX RATING 2] ... Sentence: "[EX N]" Given the above key, the readability of the sentence is (scale=1-6): [EX RATING N] Sentence: "[SENTENCE]" Given the above key, the readability of the sentence is (scale=1-6):
Table 10: Prompt provided to GPT4, GPT3.5, Aya23-8b, Llama2-7b, and Llama3.1-8b models to assess in-context learning readability assessment capabilities.